OpenStudy (anonymous):

Rectangular poster has an area of (2x^2 - 9x + 10), which describe the dimension of the poster?

geerky42 (geerky42):

I think you need to factor 2x² - 9x +10. Can you do it?

geerky42 (geerky42):

Ok so the rectangle has area of 2x² - 9x +10 and each multiple choice give you two factor (ax + b) etc. one of the answer will have two factors multipled together to be 2x² - 9x +10 So you need to factor 2x² - 9x +10. Do you know how to do it?

OpenStudy (anonymous):

No I don't remember how to factor it.

geerky42 (geerky42):

Ok so look at first term (2x²) it is a multiply of x and 2x, right? And take a look at signs, it's (__ - __ +__) So if you factor 2x² - 9x +10, it should look like (2x - __)(x - __) right?

OpenStudy (anonymous):

oh ok got it

geerky42 (geerky42):

So you can factor it now?

geerky42 (geerky42):

sorry, it's hard for me to explain how to factor...

jimthompson5910 (jim_thompson5910):

one way to factor is to solve 2x^2 - 9x + 10 = 0 for x then use the roots to find the factors

jimthompson5910 (jim_thompson5910):

you can solve 2x^2 - 9x + 10 = 0 for x using the quadratic formula

OpenStudy (campbell_st):

here is a simple method for factoring a quadratic in the form \[ax^2 + bx + c = 0\] multiply a and c in your question is 2 x 10 = 20 find the factors of 20 that add to -9.... hint they are both negative -4, -5 then write your binomials as (ax + factor 1)(ax + factor2) ------------------------- = 0 a so you have (2x -4)(2x - 5) -------------- = 0 2 there is a factor that can be removed from the first binomial and cancelled witht he denominator hope this helps 2

jimthompson5910 (jim_thompson5910):

using the quadratic formula, you can determine 2x^2 - 9x + 10 = 0 has the two solutions x = 5/2 or x = 2 get everything to one side (and 0 on the right side) and use the zero product property to get x - 5/2 = 0 or x - 2 = 0 2x - 5 = 0 or x - 2 = 0 (2x-5)(x-2) = 0 so this shows us that 2x^2 - 9x + 10 factors to (2x-5)(x-2)

geerky42 (geerky42):

Is this clear?? @avisos2830

OpenStudy (campbell_st):

so removing the common fractor you get 2(x -2)(2x - 5) -------------- = 0 2 leaves (x -2)(2x -5) = 0

OpenStudy (anonymous):

Thanks for the help!